# Leveraging Discrete Function Decomposability for Scientific Design

### v1

@@ A.6.2 Protein property predictors
Figure A3: Comparison of a standard EDA to DADO on protein property predictive models with $\mathbf{K=1000}$. For each of seven proteins of varying length, a, Amyloid (also Fig. 3a), b, AAV (also Fig. 3b), c, GB1 (also Fig. 3c), d, TDP-43 (also Fig. 3d), e, ynzC, f, CreiLOV, and g, Gcn4, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
For comparison, we show the same seven synthetic landscapes when DADO and the EDA can draw only $K=100$ samples each iteration (Fig. A4). DADO performs substantially better than the EDA on Amyloid and AAV still. On GB1, the EDA performs about as well with $K=100$ and $K=1000$. DADO is comparable to the EDA for $K=100$, but finds designs with much higher $f(x)$ when $K=1000$. For TDP-43 and ynzC, DADO finds designs with higher $f(x)$ than the EDA when $K=100$, whereas at $K=1000$, the two methods perform indistinguishably well. Both methods are comparable on CreiLOV for $K=100$ still, and DADO finds much better designs than EDA on Gcn4. Overall, as in Sec. A.6.1, we see that increasing $K$ helps, or at least does not hurt, both DADO and the EDA on all protein landscapes considered.
@@
Figure A4: Comparison of a standard EDA to DADO on protein property predictive models with $\mathbf{K=100}$. For each of seven proteins of varying length, a, Amyloid, b, AAV, c, GB1, d, TDP-43, e, ynzC, f, CreiLOV, and g, Gcn4, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=100$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
A.6.3 Investigation of protein property predictive model decomposability
We also sought to investigate how changing the threshold that determines the complexity of the junction tree would affect both predictive performance and optimization performance. the experiments, shown for experiments in which we draw 100 samples at each iteration. Specifically, we varied the distance threshold, $t$, for which pairs of residues were considered contacting to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger advantage over the standard EDA, but may be overly restrictive and yield a worse predictive model. We decided to do a case study of GB1 for $K=100$ samples at each iteration because in this setting, using the default threshold of $t=4.5\text{\AA }$ (Fig. A4c) leads to comparable performance between DADO and the EDA. We wanted to see if decomposing the model further would result in a function that’s easier for DADO to optimize without sacrificing accuracy. We find that the largest $t$ provides the highest holdout predictive accuracy, which is expected because the resulting decomposed model isn’t restricted. However, decreasing this distance down to $t=2.75\text{\AA }$ allows most of the predictive signal to remain while substantially reducing the complexity of the junction tree (Fig. A5a), enabling DADO to consistently converge on designs with high $f(x)$ within only 10 iterations (Fig. A5b), whereas for $t=4.5\text{\AA }$ and $t=9\text{\AA }$, DADO’s mean does not clearly converge even after 100 iterations, and its distribution remains dispersed as evidenced by wide shaded envelopes (Fig. A5c,d). Additionally, DADO definitively outperforms the decomposition-unaware EDA when $t=2.75\text{\AA }$ (Fig. A5b). When the predictive model is less decomposed, DADO is not as distinguishable from the EDA (Fig. A5c,d).
@@ A.6.3 Investigation of protein property predictive model decomposability
Figure A5: Investigation of GB1 predictive model decomposability. a, We varied the distance threshold, $t$, for which pairs of residues in protein GB1’s 3D structure are considered neighbors in the junction tree, using several values in $t\in[0\text{\AA },15\text{\AA }]$, so as to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger relative advantage over the standard EDA, but may yield a worse predictive model from imposing a more restrictive functional structure. The Pearson correlation on a 10% holdout set remains quite good down to and including $t=2.75\text{\AA }$, a point at which DADO provides a statistically significant win over a standard EDA as seen in panel b. Blue stars denote values of $t$ corresponding to the experiments in panels b, $t=2.75\text{\AA }$, c, $t=4.5\text{\AA }$ (also shown in Fig. A4c), and d, $t=9\text{\AA }$. Each method drew 100 samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds that dictated initialization of the search distribution. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.

### v2

@@ A.9.1 Synthetic functions: L=400
Figure A5: Comparison of a naive EDA to DADO on a synthetic problem with $L=400$. We created a random function, $f(x)$, with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. We used alphabet size $D=20$ and sequence length $L=400$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-value is from a two-sided paired t-test that the mean at the final iteration is different between methods, over the 20 seeds.
A.9.2 Synthetic functions: increasing D
For all of our synthetic experiments, we used an alphabet size of $D=20$, which reflects the typical alphabet for protein design problems. Design problems in other scientific domains might have larger alphabets so we also considered $D=50$ and $D=100$ here, keeping $L$ fixed at 100.
We observe that DADO finds designs with higher $f(x)$ than the naive EDA across all three alphabet sizes (Fig. A6). As one would expect, as $D$ grows and the design space grows combinatorially larger, it becomes increasingly difficult for both methods to optimize $f(x)$. In general, such problems require using a larger sampling budget ($K$) and/or more iterations. Although the performance gap between DADO and EDA seems to shrink as $D$ grows, the EDA has converged to a suboptimal region of the design space, whereas DADO still has a positive slope and a lot of sampling diversity. This suggests that were we to run more iterations, DADO would continue to improve, but the EDA would not.
[…]
Figure A6: Comparison of a naive EDA to DADO on synthetic problems with large alphabets. We created three random functions, $f(x)$, each with the same randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used sequence length $L=100$ and alphabet size a, $D=20$, b, $D=50$, and c, $D=100$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from a two-sided paired t-test that the mean at the final iteration is different between methods, over the 20 seeds.
A.9.3 Collected protein experiments
Here, we collect the four proteins shown in the main text (Fig. 3) alongside three other proteins we tested (details of why those four were chosen in Sec. A.6). First, we plot them with $\log(c-f(x))$ on the y-axis, for clarity when $f(x)$ is high (Fig. A7). For comparison, we then show the same experimental results with just $f(x)$ on the y-axis (Fig. A8). Whereas AAV, Amyloid, Gcn4, and TDP-43 have decomposition junction trees with nodes of cardinality five or less, the junction trees for GB1 and ynzC have many nodes with cardinality over 10, and CreiLOV has nodes with cardinality greater than 20. For GB1 and ynzC, DADO still outperforms the decomposition-unaware baselines, but not FDA, which also operates in the decomposed design space. We hypothesize that FDA performs well relative to DADO in these cases because as the junction tree nodes grow larger, the computed value functions become higher variance estimates, such that it can be better to weight the search distribution by $f(x)$ directly. CreiLOV has even larger nodes, potentially amplifying this effect. For visualization and comparison of junction trees across proteins, see Fig. A4. It would be interesting to test this hypothesis by implementing further variance-reduction techniques for DADO’s value functions.
@@ A.9.3 Collected protein experiments
Figure A7: Optimization performance on protein problems, plotted on a negative log scale. For each of seven proteins of varying length, a, AAV (also Fig. 3a), b, Amyloid (also Fig. 3b), c, Gcn4 (also Fig. 3c), d, TDP-43 (also Fig. 3d), e, GB1, f, ynzC, and g, CreiLOV, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $\log(c-f(x))$, averaged across results from 20 random seeds. We plot this quantity to make clear the differences between methods when $f(x)$ is large; $c$ is the largest $f(x)$ on a given plot, plus a small constant for numerical stability. P-values are from a two-sided paired t-test that the mean at the final iteration is different between methods, over the 20 seeds.
@@
[…]
Figure A8: Optimization performance on protein problems, plotted on a standard scale. For each of seven proteins of varying length, a, AAV (also Fig. 3a), b, Amyloid (also Fig. 3b), c, Gcn4 (also Fig. 3c), d, TDP-43 (also Fig. 3d), e, GB1, f, ynzC, and g, CreiLOV, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used standard EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from a two-sided paired t-test that the mean at the final iteration is different between methods, over the 20 seeds.
A.9.4 Investigation of protein property predictive model decomposability
We also sought to investigate how changing the threshold that determines the complexity of the junction tree would affect both predictive performance and optimization performance. Specifically, we varied the distance threshold, $t$, for which pairs of residues were considered contacting to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger advantage over the standard EDA, but may be overly restrictive and yield a worse predictive model. We decided to do a case study of GB1 for $K=100$ samples at each iteration because in this setting, using the default threshold of $t=4.5\text{\AA }$ leads to comparable performance between DADO and the EDA. We wanted to see if decomposing the model further would result in a function that’s easier for DADO to optimize without sacrificing accuracy. We find that the largest $t$ provides the highest holdout predictive accuracy, which is expected because the resulting decomposed model isn’t restricted. However, decreasing this distance down to $t=2.75\text{\AA }$ allows most of the predictive signal to remain while substantially reducing the complexity of the junction tree (Fig. A9a), enabling DADO to consistently converge on designs with high $f(x)$ within only 10 iterations (Fig. A9b), whereas for $t=4.5\text{\AA }$ and $t=9\text{\AA }$, DADO’s mean does not clearly converge even after 100 iterations, and its distribution remains dispersed as evidenced by wide shaded envelopes (Fig. A9c,d). Additionally, DADO definitively outperforms the decomposition-unaware EDA when $t=2.75\text{\AA }$ (Fig. A9b). When the predictive model is less decomposed, DADO is not as distinguishable from the EDA (Fig. A9c,d).
[…]
@@ A.9.4 Investigation of protein property predictive model decomposability
Figure A9: Investigation of GB1 predictive model decomposability. a, We varied the distance threshold, $t$, for which pairs of residues in protein GB1’s 3D structure are considered neighbors in the junction tree, using several values in $t\in[0\text{\AA },15\text{\AA }]$, so as to explore the tradeoff between accuracy of the model and decomposability. The lower the value of $t$, the stricter the decomposition (the smaller the cardinality of the largest meta-variable); thus lower $t$ should give DADO a larger relative advantage over the standard EDA, but may yield a worse predictive model from imposing a more restrictive functional structure. The Pearson correlation on a 10% holdout set remains quite good down to and including $t=2.75\text{\AA }$, a point at which DADO provides a statistically significant win over a standard EDA as seen in panel b. Blue stars denote values of $t$ corresponding to the experiments in panels b, $t=2.75\text{\AA }$, c, $t=4.5\text{\AA }$, and d, $t=9\text{\AA }$. Each method drew 100 samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds that dictated initialization of the search distribution. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
A.9.5 Decomposition graph robustness
How crucial is it to guess the decomposition perfectly a priori? One way to investigate this is to mutate the decomposition graph and observe how the resulting decomposed predictive models’ predictive accuracies change. In our first set of experiments, we varied the contact threshold ($t$) used on the AlphaFold 3D structure to determine connectivity; lowering $t$ will gradually remove more distant contacts, whereas increasing $t$ will gradually add more distant contacts. Given different decomposition graphs, we then trained a decomposed predictive model for each. For our second set of experiments, we performed random mutations to the t=4.5t=4.5Å  decomposition graph (used in Sec. 4) to study how robust prediction is to missing / extra edges. In particular, for each of $N\in[-50,-10,-5,-1,1,5,10,50]$, we randomly sampled $N$ edges to remove/add. We implemented a check to ensure that the graph doesn’t become disconnected, so for some experiments, we cut off $N$ at the largest number of edges that could be removed resulting in a chain graph. We repeated this procedure 10 times.
In both of our experiments, we considered holdout accuracy as measured by the Pearson correlation coefficient between assay labels and model predictions. All decomposed models for each protein used the same hyperparameters that were chosen via 5-fold cross validation on the full dataset for the base t=4.5t=4.5Å  decomposed model, for expediency. We also compared to a neural network without any decomposition (i.e., an all-edges model), denoted “Naive NN” in our plots. For this model we performed an additional 5-fold cross validation on the full dataset for each protein.
We observe that the holdout accuracy of the non-decomposed predictive model and the decomposed models of varying $t$ all tend to fall within a relatively small range (Fig. A10), suggesting that using our decomposed predictive model does not constitute a substantial sacrifice compared to a full (naive) model, and that our decomposition is somewhat robust to changing the contact threshold. When we randomly mutate the decomposition graph, we find that holdout accuracy is generally concentrated within an even smaller range, with some exceptions when many edges are added or removed (Fig. A11). Interestingly, for Gcn4, the decomposed predictive models outperform the naive model, though it’s worth noting that this prediction task is especially hard due to the relatively uniform dispersion of sequences throughout the design space, and the small size of the dataset. Both models fit the training data well (not shown here).
[…]
Figure A10: Investigation of predictive model decomposability by contact threshold. We varied the contact threshold, $t$, for which pairs of residues in each protein’s 3D structure are considered neighbors in the decomposition graph, using several values in $t\in[0\text{\AA },15\text{\AA }]$, so as to explore the tradeoff between accuracy of the model and decomposability. We studied a, AAV, b, Amyloid, c, Gcn4, d, TDP-43, and e, GB1.
[…]
Figure A11: Investigation of predictive model decomposability by random graph mutation. For each of a, AAV, b, Amyloid, c, Gcn4, d, TDP-43, and e, GB1, we randomly added or removed edges from the t=4.5t=4.5Å  decomposition graph. In particular, we randomly sampled $N\in[-50,-10,-5,-1,1,5,10,50]$ edges to remove/add. For certain experiments we limit NN to avoid creating disconnected graphs. We repeated this procedure 10 times.
A.9.6 Decomposition graph robustness: bottom 50% data
Herein, we repeated the same analyses as in Sec. A.9.5, except using the bottom half of the training set (i.e., the datapoints with the lowest assay labels). We used the same exact holdout sets as in Sec. A.9.5, such that decomposed predictive models are tested against sequences with both high and low assay labels. We used the same hyperparameters for all of the predictive models before, which were cross-validated on the full dataset.
Overall, we observe that compared to Fig. A10 and Fig. A11, holdout accuracy as measured by the Pearson correlation coefficient between assay labels and model predictions is lower for all models and all proteins (Fig. A12, Fig. A13). Interestingly, for several proteins, the more decomposed predictive models (lower $t$) performed better than the less decomposed models (higher $t$), as well as the full (naive) model (Fig. A12). This makes sense, as more complex models are more prone to overfitting when the training data is shifted from the holdout distribution, and as the training dataset gets smaller. This trend is also reflected in the random graph mutation experiments (Fig. A13). Generally, the same trends observed in Sec. A.9.5 hold here, such as predictive accuracy being relatively robust to changing $t$ and to randomly mutating the graph (up to a point).
[…]
Figure A12: Investigation of predictive model decomposability by contact threshold (bottom 50% data). We varied the contact threshold, $t$, for which pairs of residues in each protein’s 3D structure are considered neighbors in the decomposition graph, using several values in $t\in[0\text{\AA },15\text{\AA }]$, so as to explore the tradeoff between accuracy of the model and decomposability. We studied a, AAV, b, Amyloid, c, Gcn4, d, TDP-43, and e, GB1. We used the same holdout set as in Fig. A10, but only trained on the bottom 50% of data by assay label.
[…]
Figure A13: Investigation of predictive model decomposability by random graph mutation (bottom 50% data). For each of a, AAV, b, Amyloid, c, Gcn4, d, TDP-43, and e, GB1, we randomly added or removed edges from the t=4.5t=4.5Å  decomposition graph. In particular, we randomly sampled $N\in[-50,-10,-5,-1,1,5,10,50]$ edges to remove/add. For certain experiments we limit NN to avoid creating disconnected graphs. We repeated this procedure 10 times. We used the same holdout set as in Fig. A11, but only trained on the bottom 50% of data by assay label.
A.9.7 Runtime analysis
We measure the wall-clock time it takes to run a single distributional optimization algorithm for both DADO and the naive EDA (Table 1). We fix both to run for 100 iterations, drawing 100 samples at each, using the same architectures as described in Sec. A.8.1, and on a single GPU. We report times just to give a rough sense of runtime and scaling; one could definitely further optimize our code.
Table 1: Runtime comparison of EDA and DADO, varying $L$ and $D$.
Problem Length (L) Alphabet size (D) EDA runtime DADO runtime
Synthetic tree 25 20 2.8 min. 2.9 min.
Synthetic tree 50 20 6.3 min. 7.0 min.
Synthetic tree 100 20 16.3 min. 16.1 min.
Synthetic tree 200 20 42.7 min. 36.2 min.
Synthetic tree 400 20 139.5 min. 114.2 min.
Synthetic tree 100 50 14.0 min. 16.0 min.
Synthetic tree 100 100 14.5 min. 18.3 min.
AAV2 capsid protein 28 20 2.1 min. 3.4 min.
TDP-43 protein 84 21 6.7 min. 10.5 min.
A few trends stand out. First, the EDA and DADO take roughly the same amount of time to run, with DADO being faster sometimes. This speed-up may come from using a factorized search distribution, which requires only conditioning on parent variables as opposed to all preceding variables (autoregressive). In general, problems with larger $L$ take longer to run, and this scaling is worse than linear for our implementation. We also notice that when increasing $D$ and holding $L$ fixed at 100, the runtime is similar. A single run of either DADO or EDA does not generally require the full memory of a 16GB GPU, so in practice, we’re able to parallelize runs within GPUs and across multiple GPUs in order to perform our hyperparameter sweep and random seed replicates efficiently.